OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30

OpenStudy (campbell_st):

for x-intercepts let y = 0 so you need to solve 15x - 0 = 30 for y-intercepts let x = 0 so you need to solve 0 - 5x = 30 this will give the intercepts

OpenStudy (anonymous):

x-intercept = (2, 0); y-intercept = (0, -6)

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

correct